# Equilibrium under Time-Inconsistency: A New Existence Theory by Vanishing Entropy Regularization

### 11 Mar 2026 v1

@@ Abstract
This paper studies time-inconsistent stochastic control problems in a continuous-time setting. The time-inconsistency arises from the initial-time dependence such as the non-exponential discounting. The classical approach typically relates the existence of equilibrium to the classical solution of the equilibrium HJB equation (EHJB), whose existence is still an open problem under general model assumptions. We resolve this challenge by invoking the entropy regularization on relaxed policies. First, by employing some fixed point arguments, we establish the existence of classical solution of the exploratory equilibrium HJB equation (EEHJB) and obtain the Gibbs form characterization of the regularized equilibrium. Next, we develop some delicate PDE estimates for the solution of the EEHJB and its derivatives, facilitating the novel convergence analysis under suitable norms as the entropy regularization vanishes. It is shown that the solutions of EEHJB converges to a weak solution of a generalized EHJB, which is associated with the limit of the regularized equilibrium. This core convergence result allows us to conduct the verification to conclude that the limiting regularized equilibrium constitutes the equilibrium in the original problem. We thus contribute a new sufficient condition to the literature for the equilibrium in diffusion models under time-inconsistency without resorting to strong regularity assumptions of the EHJB.
Keywords: Time-inconsistent control, entropy regularization, exploratory equilibrium HJB equation, convergence analysis, generalized equilibrium HJB equation, relaxed equilibrium
@@ 1 Introduction
Vast real-life financial and economic problems are compromised to be time-inconsistent due to the non-exponential discounting adopted by the decision maker, that is, a policy deemed optimal today may no longer remain optimal at future dates. In response to the failure of global optimality, a more suitable alternative solution, initiated by [24], is to look for the subgame perfect Nash equilibrium for the intra-personal game among the decision maker’s current and future selves. In the continuous-time setting, the seminal study of [2] considered a perturbation of the control for an infinitesimal period of time and used the limit of time-discretization to derive the extended HJB equation as a system of nonlinear PDEs. By proving a rigorous verification theorem, the equilibrium can be characterized by the classical solution of the extended HJB equation in [2]. From the approximation perspective by discrete-time intra-personal games, [31] formulated a nonlinear and nonlocal PDE, called the equilibrium HJB equation (EHJB), which is mathematically equivalent to the extended HJB system in [2] after some transformations. These two pioneer studies inspired a wave of interesting work on continuous-time time-inconsistent control problems, primarily focusing on the analysis of extended HJB system or EHJB in different models. As the verification theorem plays a vital role in establishing the existence and characterization of equilibrium, the conventional approach in the literature stands on the stringent regularity conditions of the classical solution to the extended HJB system or EHJB. However, analyzing the existence of classical solution to the general system of nonlinear and nonlocal PDEs remains an challenging open problem. Recently, [20] investigated the short-time classical solution to the general EHJB under specific model assumptions and [21] further established the existence of global solution of EHJB for time-inconsistent game problems relying on stronger conditions on model coefficients and cost functions. When these restrictive model assumptions are violated or when the existence of classical solution to extended HJB or EHJB is unknown, equilibrium in continuous-time models under time-inconsistency remains nearly unexplored.
@@
Apart from the scope of RL, it has also been unfolded in some recent studies that the explicit Gibbs measure characterization of the optimal relaxed control under entropy regularization bears some mathematical simplifications in exercising fixed point arguments. By further establishing the convergence as entropy regularization vanishes, it offers a new route to confirm the existence of solution to the original problem as the limit from the regularized problems. Such an approach was initiated in [1], which studied time-inconsistent MDP problems under general discounting and established the existence of relaxed equilibrium in both discrete-time and continuous-time settings using the approach of vanishing entropy regularization; Subsequently, [32] investigated the time-inconsistent mean-field MDP of stopping and proved that the regularized equilibrium converges to the relaxed equilibrium; In discrete-time major-minor mean-field games of stopping, [33] employed the vanishing entropy regularization approach in the major-player’s problem and proved the existence of mean-field equilibrium; [8] studied the convergence of the solutions to the exploratory extended HJB equations for continuous-time mean-variance stopping problems by assuming the existence of classical solutions, which leads to the existence of equilibrium in the original time-inconsistent stopping problem.
Inspired by some observed technical convenience of entropy regularization in aforementioned studies, this paper aims to revisit the time-inconsistent control problems when the drift coefficient of state dynamics is controlled and develop a new theory for the existence of equilibrium by analyzing the convergence of PDEs under vanishing entropy regularization. On one hand, unlike the discrete-time models in [1], [32], [33], we really have to shift the focus from the convergence of operators to the more delicate convergence analysis of associated HJB equations arising from the continuous-time setting. Some technical estimations of the solution and its derivatives are inevitable. On the other hand, in sharp contrast to [25] and [11] for classical time-consistent optimal control problems, we can no longer resort to some stability results of viscosity solutions in the literature. Indeed, the dynamic programming principle fails in our model and we are in the absence of necessary condition as viscosity characterization of the equilibrium value function. The rigorous convergence analysis and the verification of equilibrium in our general setting calls for some novel technical tools and arguments. Comparing with the recent work [8], we do not assume some priory LQ solution structure and do not impose the strong assumption of existence of classical solution to the EHJB for the proof of convergence when entropy regularization vanishes.
@@
Firstly, we consider the time-inconsistent stochastic control problem under the Shannon entropy regularization and derive the exploratory equilibrium HJB equation (EEHJB) system (3.5)–(3.6). Our goal in this step is to establish the existence of classical solution by invoking the fixed point argument on a specialized compact space with the help of the Gibbs-form policy operator in (3.7) for the equilibrium. Thanks to the sublinear growth estimate of the entropy term in Lemma 3.1 for small weighting parameter $\lambda$, we are able to first establish some Hölder norm estimates for the Gibbs-form operator $\pi(x,a):=\Gamma_{\lambda}(x,D_{x}w(x),a)$ with an initial input $w(x)\in{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})$ under certain model assumptions (see Lemma 3.2). We then further show its resulting value function $V_{\lambda}^{\pi}$ satisfies the PDE in the EEHJB system with some weak type of Hölder estimates (see Lemma 3.3). Building upon these estimates, the equilibrium essentially boils down to the fixed point of the operator $\Phi_{\lambda}(w):=V^{\Gamma_{\lambda}(D_{x}w(0,x))}$, for which a key step is to identify a suitable compact set. Our first main technical contributions is to propose to work with $\mathcal{M}_{\lambda}$ defined by (3.30) as a specialized compact subset of ${\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\beta/2,\beta}(\mathbb{T}\times\mathbb{R}^{d})$ together with a tailor-made weighted global Hölder norm, which ensures that $\Phi_{\lambda}$ is a continuous mapping from $\mathcal{M}_{\lambda}$ to itself after some careful arguments (see Lemma 3.4). Thanks to these preparations, Schauder fixed-point theorem can be called such that the value function under the fixed point satisfies the EEHJB system with the desired regularity (see Theorem 3.1), which enables us to conclude the existence of regularized equilibrium by the verification arguments in Proposition 3.1.
@@
Secondly, we cope with the more challenging task by analyzing the convergence of EEHJB system towards the original EHJB system when the entropy parameter $\lambda{\to}0$. To this end, we consider a pair of sequence $(v^{n},\pi^{n})$ satisfying the nonlinear PDEs (4.1) and (4.2) in the EEHJB system, respectively. Employing some norm estimates and diagonal arguments, we first show the existence of a subsequence $v^{n_{k}}$ converging locally on $D_{N}$ to a function $v^{\infty}\in{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})\cap W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})$ under the Hölder norm ${\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{N})$ for each $N\in\mathbb{N}$, whose derivatives also converge to the derivative of $v^{\infty}$ in the distribution sense on each local domain $D_{N}$ (see item (i) of Lemma 4.1). For the same subsequence, we can prove the weak convergence of $\pi^{n_{k}}$ towards a Borel measurable $\pi^{\infty}$ with the aid of Young measure theory. As the second main theoretical contribution, we address the most intricate step to relate the limit $v^{\infty}$ to the solution of the PDE under the limit $\pi^{\infty}$ such that $v^{\infty}=V^{\pi^{\infty}}$ by developing some novel verification arguments for $v^{\infty}$ to be the solution of a limiting PDE based on the convergence in distribution and the delicate treatment to extend $v^{\infty}$ from the local limit on $D_{N}$ to the entire domain (see item (iii) of Lemma 4.1). Our convergence result can be interpreted as the stability result of the solution to the EEHJB system with respect to $\lambda{\to}0$ in the time-inconsistent setting, the first kind of result in the existing literature.
@@
Furthermore, to complete the final goal of showing that the limit $\pi^{\infty}$ constitutes a relaxed equilibrium in the original problem, we highlight another new technical contribution as our verification arguments. Indeed, after extending from the local domain to the entire domain using the converging sequence $v^{n_{k}}$, it is revealed that the limit $v^{\infty}\in{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})\cap W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})$ only satisfies the EHJB system in a distributional sense. However, thanks to the established convergence of $v^{n_{k}}$ to $v^{\infty}$ and each $v^{n_{k}}$ corresponds to the classical solution of EEHJB, we can take full advantage of convergence in distribution and Itô-Krylov formula, together with some delicate localization arguments and estimations around $t=0$, to conclude that $\pi^{\infty}$ indeed fulfils the definition of relaxed equilibrium in (2.4); see Theorem 4.1. Unlike most existing studies, our verification proof does not lean on the classical solution. As a byproduct of our verification arguments, as stated in Corollary 4.1, we establish an interesting weaker sufficient condition, new to the literature, for the characterization of equilibrium, which only requires the solution $u\in{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})\cap W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})$ to satisfy the inequality (4.17) a.e. for $(t,x)\in[0,\epsilon_{0}]\times\mathbb{R}^{d}$, regarded as a weak-type EHJB, in the distributional sense. Our new existence result, stemming from the convergence analysis of EEHJB, may open a new venue to study time-inconsistent control problems when the strong regularity conditions of EHJB are difficult to prove or are completely absent.
@@
Meanwhile, it is also worth mentioning a recent study [13] that established the policy iteration convergence for time-inconsistent control problems under entropy regularization, which partially laid the theoretical foundation for the design of some policy-iteration-based RL algorithms in time-inconsistent settings. Similar to the time-consistent counterpart, one natural question arises: although for the fixed entropy regularization, the greedy iteration of the Gibbs-form policies can lead to the regularized equilibrium, do we know that the regularized equilibrium can be arbitrarily close to the equilibrium in the original problem when the entropy term vanishes? In other words, can the learned solution in the exploratory formulation approximate the true equilibrium when the temperature parameter is set to be small? The main result of this paper addresses this question in the model with drift control and confirms that the regularized equilibrium can effectively converge to a relaxed equilibrium in the original problem, which can justify the rationale behind the choice of small temperature parameters in RL algorithms even in some time-inconsistent settings. Moreover, comparing with [13] that contributed to the existence of classical solution to EEHJB using the constructive proof of policy iteration, we contribute a fundamentally distinct proof for the existence of classical solution to EEHJB using the more direct fixed point arguments, in which some newly developed PDE estimates also play important roles in the latter part of convergence analysis when entropy regularization vanishes.
@@
The rest of the paper is organized as follows. Section 2 introduces the problem formulation of time-inconsistent control problems, the definition of relaxed equilibrium and its associated EHJB system. Section 3 formulates the problem under entropy regularization and the induced EEHJB system. For small entropy parameter $\lambda$, the existence of classical solution to the EEHJB system is established therein that characterizes the regularized equilibrium in Gibbs-form. Section 4 further studies the convergence analysis of the solutions of EEHJB towards a weak solution of the EHJB in distributional sense and developed new verification arguments to confirm that the limit of regularized equilibrium is the equilibrium in the original time-inconsistent control problem when entropy regularization vanishes.
@@ 1.1 Notations
Let $\mathbb{N}$ be the set of all positive integers and denote the infinite time horizon as $\mathbb{T}:=[0,\infty)$. Given the $m$-dimensional Euclidean space $\mathbb{R}^{m}$, denote by $|\cdot|$ the Euclidean norm, by $\operatorname{\texttt{Leb}}(\cdot)$ the Lebesgue measure, and by $B_{r}(x)$ the ball centered at $x$ with radius $r$.
Given $w(t,x):\mathbb{T}\times{\mathbb{R}^{d}}\mapsto{\mathbb{R}}$, let $\partial_{t}f$ denote the right derivative on the time variable $t$ and $D_{x}w$ (resp. $D^{2}_{x}w$) denote the first (resp. second) partial derivative vector (Hessian) on the space variable $x$. We first recall notations of Hölder spaces. Given a simple connected domain $D\subset\mathbb{T}\times{\mathbb{R}^{d}}$, define
$\|w\|_{{\mathcal{C}}^{0}(D)}$= $\sup_{(t,x)\in D}|w(t,x)|=\|w\|_{L^{\infty}(D)},\quad[w]_{{\mathcal{C}}_{\alpha/2,\alpha}(D)}$= $\sup_{(t,x)\neq(s,y)\in D}\frac{|w(t,x)-w(s,y)|}{(|t-s|+|x-y|^{2})^{\alpha/2}},$
@@
$\|w\|_{{\mathcal{C}}^{l,k}_{\alpha/2,\alpha}(D)}:=\sum_{0\leq l\leq 1,0\leq|\alpha|_{l_{1}}\leq k}\|\partial^{l}_{t}D^{a}_{x}w\|_{{\mathcal{C}}_{\alpha/2,\alpha}(D)}<\infty.$
@@
and we denote by ${\mathcal{C}}^{1,k}_{\alpha/2,\alpha}(D)$ (resp. ${\mathcal{C}}^{1,k,\operatorname{\textbf{wg}}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$) the normed space of all functions with finite norm $\|w\|_{{\mathcal{C}}^{1,k}_{\alpha/2,\alpha}(D)}$ (resp. finite norm $\|\cdot\|_{{\mathcal{C}}^{1,k,\operatorname{\textbf{wg}}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}$).
We also apply the Sobolev norm $\|w\|_{W^{1,k}_{p}(D)}:=\sum_{0\leq l\leq 1,0\leq|\alpha|_{l_{1}}\leq k}\|\partial^{l}_{t}D^{a}_{x}w\|_{L^{p}(D)}$, and we define the uniformly local Sobolev norm
$\|w\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})}$= $\sum_{0\leq l\leq 1,0\leq|\alpha|_{l_{1}}\leq k}\left(\sup_{(t,x)\in\mathbb{T}\times\mathbb{R}^{d}}\|\partial^{l}_{t}D^{a}_{x}w\|_{L^{p}(D_{1}(t,x))}\right)$
In the present paper, let $0<\alpha<1$ be an arbitrarily fixed Hölder constant, and we shall let the power $p$ in any Sobolev norm equal to $\frac{d+2}{1-\alpha}$. However, the whole arguments are still valid for other Hölder constants and other possible powers.

### 04 Apr 2026 v2

@@ Abstract
This paper develops a framework for establishing the existence of solutions to the equilibrium Hamilton–Jacobi–Bellman (EHJB) equation arising in time-inconsistent stochastic control problems. The time-inconsistency in our setting arises from the initial-time dependence such as the non-exponential discounting. The classical approach typically relates the existence of equilibrium to the classical solution of the EHJB, whose existence is still an open problem under general model assumptions. We resolve this challenge by building on a vanishing entropy regularization approach. Using fixed-point arguments, we first establish the existence of classical solutions to the exploratory equilibrium Hamilton–Jacobi–Bellman Equation (EEHJB) by deriving a series of delicate PDE estimates for the solution and its derivatives. Building on these estimates for the solution of the EEHJB and its derivatives, we then conduct a rigorous convergence analysis under suitable norms as the entropy regularization vanishes. Our main result shows that solutions of the EEHJB converge to a strong solution of the original EHJB, corresponding to the limit of the regularized equilibria. This convergence yields a verification argument ensuring that the limiting relaxed equilibrium indeed constitutes an equilibrium for the original time-inconsistent control problem. We thus establish the well-posedness of the EHJB and the existence of equilibria in diffusion models under time-inconsistency, without resorting to conventional stringent regularity assumptions of the EHJB.
Keywords: time-inconsistent stochastic control, entropy regularization, exploratory equilibrium HJB equation, convergence analysis, generalized equilibrium HJB equation, relaxed equilibrium
@@ 1 Introduction
Vast real-life financial and economic problems are compromised to be time-inconsistent due to the non-exponential discounting adopted by the decision maker, that is, a policy deemed optimal today may no longer remain optimal at future dates. In response to the failure of global optimality, a more suitable alternative solution, initiated by [24], is to look for the subgame perfect Nash equilibrium for the intra-personal game among the decision maker’s current and future selves. In the continuous-time setting, the seminal study of [2] considered a perturbation of the control for an infinitesimal period of time and used the limit of time-discretization to derive the extended HJB equation as a system of nonlinear PDEs. By proving a rigorous verification theorem, the equilibrium can be characterized by the classical solution of the extended HJB equation in [2]. From the approximation perspective by discrete-time intra-personal games, [31] formulated a nonlinear and nonlocal PDE, called the equilibrium HJB equation (EHJB), which is mathematically equivalent to the extended HJB system in [2] after some transformations. These two pioneer studies inspired a wave of interesting work on continuous-time time-inconsistent control problems, primarily focusing on the analysis of extended HJB system or EHJB in different models. As the verification theorem plays a vital role in establishing the existence and characterization of equilibrium, the conventional approach in the literature stands on the stringent regularity conditions of the classical solution to the extended HJB system or EHJB. However, analyzing the existence of classical solution to the general system of nonlinear and nonlocal PDEs remains a challenging open problem. Recently, [20] investigated the short-time classical solution to the general EHJB under specific model assumptions and [21] further established the existence of global solution of EHJB for time-inconsistent game problems relying on stronger conditions on model coefficients and cost functions. When these restrictive model assumptions are violated or when the existence of classical solution to extended HJB or EHJB is unknown, equilibrium in continuous-time models under time-inconsistency remains nearly unexplored.
@@
Apart from the scope of RL, it has also been unfolded in some recent studies that the Gibbs measure characterization of the optimal relaxed control under entropy regularization brings mathematical simplifications in exercising fixed point arguments. By further establishing the convergence as entropy regularization vanishes, it offers a new route to confirm the existence of solution in the original problem as the limit from the regularized problems. Such an approach was initiated in [1], which studied time-inconsistent MDP problems under general discounting and established the existence of relaxed equilibrium in both discrete-time and continuous-time settings using the approach of vanishing entropy regularization; Subsequently, [32] investigated the time-inconsistent mean-field MDP of stopping and proved that the regularized equilibrium converges to the relaxed equilibrium; In discrete-time major-minor mean-field games of stopping, [33] employed the vanishing entropy regularization approach in the major-player’s problem and proved the existence of mean-field equilibrium; [8] studied the convergence of the exploratory extended HJB equations for continuous-time mean-variance stopping problems by assuming the existence of classical solutions, which leads to the existence of equilibrium in the original problem.
Inspired by some observed technical convenience of entropy regularization in aforementioned studies, this paper aims to revisit the time-inconsistent control problems when the drift coefficient of state dynamics is controlled and develop a new theory for the existence of equilibrium by analyzing the convergence of PDEs under vanishing entropy regularization. On one hand, unlike the discrete-time models in [1], [32], [33], we have to shift the focus from the convergence of operators to the more delicate convergence analysis of associated HJB equations arising from the continuous-time setting. Some technical estimations of the solution and its derivatives are inevitable. On the other hand, in sharp contrast to [25] and [11] for classical time-consistent optimal control problems, we can no longer resort to some stability results of viscosity solutions in the literature. Indeed, the dynamic programming principle fails in our model and we are in the absence of necessary viscosity characterization of the equilibrium value function. The rigorous convergence analysis and the verification of equilibrium in our general setting call for new technical tools and PDE analysis. Comparing with the recent work [8], we do not assume any priory LQ solution structure and do not impose the restrictive assumption on the existence of classical solution to the EHJB for the proof of convergence analysis.
@@
Firstly, we consider the time-inconsistent stochastic control problem under the Shannon entropy regularization and derive the exploratory equilibrium HJB equation (EEHJB) system (3.5)–(3.6). The main contribution of this step is establishing the existence of classical solution to the EEHJB system by invoking the fixed point argument on a specialized compact space with the help of the Gibbs-form policy operator in (3.7). Thanks to the sublinear growth estimate of the entropy term in Lemma 3.1, we first derive Hölder norm estimates for the Gibbs-form operator $\pi(x,a):=\Gamma_{\lambda}(x,D_{x}w(x),a)$ with an initial input $w(x)\in{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})$ (see Lemma 3.2). We then show that the associated value function $V_{\lambda}^{\pi}$ satisfies the EEHJB system and enjoys certain weak type of Hölder regularity estimates (see Lemma 3.3). Building upon these estimates, the equilibrium essentially boils down to the fixed point of the operator $\Phi_{\lambda}(w):=V^{\Gamma_{\lambda}(D_{x}w(0,x))}$, for which a key step is to identify a suitable compact set. Our first main contribution is to propose to work with $\mathcal{M}_{\lambda}$ defined by (3.33) as a specialized compact subset of ${\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\beta/2,\beta}(\mathbb{T}\times\mathbb{R}^{d})$ together with a tailor-made weighted global Hölder norm. This construction ensures that $\Phi_{\lambda}$ defines a continuous self-mapping on $\mathcal{M}_{\lambda}$ (see Lemma 3.5). Thanks to these preparations, Schauder fixed-point theorem can be called such that the value function under the fixed point satisfies the EEHJB system with the desired regularity (see Theorem 3.1), which enables us to conclude the existence of regularized equilibrium by some standard verification arguments.
@@
Secondly, we cope with the more challenging task by analyzing the convergence of EEHJB system towards the original EHJB system when the entropy parameter $\lambda{\to}0$. To this end, we consider a sequence of pairs $(v^{n},\pi^{n})_{n\in\mathbb{N}}$ satisfying the nonlinear PDEs (4.3) and (4.4) in the EEHJB system, respectively. Building on the norm estimates established in the previous section and employing a diagonal argument, we first show the existence of a subsequence $v^{n_{k}}$ converging locally on $D_{N}$ to a function $v^{\infty}\in{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})\cap W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})$ under the Hölder norm ${\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{N})$ for each $N\in\mathbb{N}$, whose derivatives also converge to the derivative of $v^{\infty}$ in the distribution sense on each local domain $D_{N}$ (see item (i) of Lemma 4.1). Along the same subsequence, we establish the weak convergence of $\pi^{n_{k}}$ towards a Borel measurable $\pi^{\infty}$ with the aid of Young measure theory. As the second main contribution, we address the most intricate step to relate the limit $v^{\infty}$ to a strong solution to the EHJB equation in the original model under the limit $\pi^{\infty}$ such that $v^{\infty}=V^{\pi^{\infty}}$, see Theorem 4.2. We achieve this core convergence result by developing novel arguments to show that $v^{\infty}$ is locally a strong solution of the EHJB and $v^{\infty}$ can be extended from the local limit on $D_{N}$ to the entire domain (see item (iii) of Lemma 4.1). As a consequence of this convergence result, we establish the existence of a strong solution to the nonlinear EHJB system in Theorem 4.2 for the original time-inconsistent control problem, a first kind of result in the literature. Moreover, our convergence analysis can also be interpreted as the stability result of the solution to the EEHJB system with respect to $\lambda{\to}0$ in the time-inconsistent setting, which has not been explored in the literature.
@@
Furthermore, our final goal is to show that the limit $\pi^{\infty}$ constitutes a relaxed equilibrium in the original problem, for which we highlight a new technical contribution: verification arguments based on the previous convergence analysis. In fact, the limit $v^{\infty}\in{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})\cap W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})$ is identified as a strong, rather than classical, solution to the original EHJB system. However, thanks to the established convergence of $v^{n_{k}}$ to $v^{\infty}$ and the fact that each $v^{n_{k}}$ is a classical solution of the EEHJB, we take full advantage of the convergence and the Itô-Krylov formula, together with localization arguments and estimates around $t=0$, to conclude that $\pi^{\infty}$ indeed satisfies the definition of relaxed equilibrium in (2.4) (see Theorem 4.1). Unlike most existing studies, our verification proof does not lean on the classical solution. Instead, we establish a new and weaker sufficient condition in Theorem 4.1 for the existence and characterization of equilibrium, which only requires a strong solution in ${\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})\cap W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})$ to the original EHJB. Notably, Theorem 4.2 guarantees via a convergence argument that such a strong solution to the EHJB system always exists, thereby establishing the general existence of equilibrium. Therefore, our new existence theory (Theorem 4.1 and Theorem 4.2 ) opens a new avenue to study time-inconsistent control problems when the strong regularity conditions (such as classical solution) of the EHJB are difficult to prove or simply unavailable.
@@
Meanwhile, it is worth mentioning a recent study [13] that established the policy iteration convergence for time-inconsistent control problems under entropy regularization, which partially laid the theoretical foundation for the design of some policy-iteration-based RL algorithms in time-inconsistent settings. Similar to the time-consistent counterpart, one natural question arises: although for a fixed entropy regularization, the greedy iteration of the Gibbs-form policies can lead to the regularized equilibrium, can the learned solution in the exploratory formulation approximate the true equilibrium when the temperature parameter is set to be small? As a byproduct of our main results, we provide an affirmative answer to this question by showing that the regularized equilibrium converges to a relaxed equilibrium of the original problem. This finding offers theoretical justification for the use of small temperature parameters in reinforcement learning algorithms, even in time-inconsistent settings. Moreover, in contrast to [13], where the existence of classical solutions to the EEHJB equation is established via a constructive policy iteration argument restricted to the finite-horizon setting, we develop a fundamentally different approach based on direct fixed-point arguments. In addition, the new PDE estimates developed in our analysis play a crucial role in the subsequent convergence study as the entropy regularization vanishes.
@@
The rest of the paper is organized as follows. Section 2 introduces the formulation of time-inconsistent control problems, the definition of relaxed equilibrium and the associated EHJB system. Section 3 formulates the problem under entropy regularization and derives the induced EEHJB system. The existence of a classical solution to the EEHJB system is established therein, together with carefully tailored norm estimates that play a crucial role in the subsequent convergence analysis. Section 4 carries out the convergence analysis of the solutions to the EEHJB towards a strong solution of the EHJB, and developed new verification arguments to confirm that the limit of regularized equilibrium is the equilibrium in the original time-inconsistent control problem as entropy regularization vanishes.
@@ 1.1 Notations
Let $\mathbb{N}$ be the set of all positive integers and denote the infinite time horizon as $\mathbb{T}:=[0,\infty)$. In the $m$-dimensional Euclidean space $\mathbb{R}^{m}$, we denote by $|\cdot|$ the Euclidean norm, by $\operatorname{\texttt{Leb}}(\cdot)$ the Lebesgue measure, and by $B_{r}(x)$ the ball centered at $x$ with radius $r$.
Given $w(t,x):\mathbb{T}\times{\mathbb{R}^{d}}\to{\mathbb{R}}$, let $\partial_{t}f$ denote the right derivative on the time variable $t$ and $D_{x}w$ (resp. $D^{2}_{x}w$) denote the first (resp. second) partial derivative vector (Hessian) on the space variable $x$. We first introduce the notation for Hölder spaces. Given a simple connected domain $D\subset\mathbb{T}\times{\mathbb{R}^{d}}$, define
$\|w\|_{{\mathcal{C}}^{0}(D)}=\sup_{(t,x)\in D}|w(t,x)|=\|w\|_{L^{\infty}(D)},\quad[w]_{{\mathcal{C}}_{\alpha/2,\alpha}(D)}=\sup_{\begin{subarray}{c}(t,x)\neq(s,y)\in D\\ (|t-s|+|x-y|^{2})\leq 1\end{subarray}}\frac{|w(t,x)-w(s,y)|}{(|t-s|+|x-y|^{2})^{\alpha/2}},$
@@
$\|w\|_{{\mathcal{C}}^{l,k}_{\alpha/2,\alpha}(D)}:=\sum_{0\leq j\leq l,0\leq|a|_{l_{1}}\leq k}\|\partial^{j}_{t}D^{a}_{x}w\|_{{\mathcal{C}}_{\alpha/2,\alpha}(D)}<\infty.$
The Hölder seminorms are also applied when restricted to either the spatial or time variable alone, e.g., $f(\cdot)]_{\alpha(O)}=\sup_{\begin{subarray}{c}x\neq y\in O\\ |x-y|\leq 1\end{subarray}}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}$ for $f$ defined on $O\subset\mathbb{R}^{d}$, and $g(\cdot)]_{\alpha/2(O)}=\sup_{\begin{subarray}{c}t\neq s\in O\\ |t-s|\leq 1\end{subarray}}\frac{|g(t)-g(s)|}{|t-s|^{\alpha/2}}$ for $g$ defined on $O\subset\mathbb{T}$.
@@
and we denote by ${\mathcal{C}}^{1,k}_{\alpha/2,\alpha}(D)$ (resp. ${\mathcal{C}}^{1,k,\operatorname{\textbf{wg}}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$) the normed space of all functions with finite norm $\|\cdot\|_{{\mathcal{C}}^{1,k}_{\alpha/2,\alpha}(D)}$ (resp. finite norm $\|\cdot\|_{{\mathcal{C}}^{1,k,\operatorname{\textbf{wg}}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}$).
We also use the Sobolev norm $\|w\|_{W^{1,k}_{p}(D)}:=\sum_{0\leq l\leq 1,0\leq|a|_{l_{1}}\leq k}\|\partial^{l}_{t}D^{a}_{x}w\|_{L^{p}(D)}$, and we further define the uniformly local Sobolev norm on the whole space $\mathbb{T}\times\mathbb{R}^{d}$
$\|w\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})}$= $\sum_{0\leq l\leq 1,0\leq|a|_{l_{1}}\leq k}\left(\sup_{(t,x)\in\mathbb{T}\times\mathbb{R}^{d}}\|\partial^{l}_{t}D^{a}_{x}w\|_{L^{p}(D_{1}(t,x))}\right)$
Throughout the paper, let $0<\alpha<1$ be an arbitrarily fixed Hölder constant, and we shall let the power $p$ in any Sobolev norm equal to $\frac{d+2}{1-\alpha}$. However, the whole arguments are still valid for other Hölder constants and other possible powers.